# LLM-Driven Multi-step Translation from C to Rust using Static Analysis

### v2

@@ Appendix A A Non-Exhaustive List of Differences Between C and Rust
Appendix B Algorithm for Task Division
@@
3:$L_{sorted}$: List of groups resolving dependencies
4:$L_{sorted}\leftarrow\emptyset$triangleright Empty list
@@
9:         $L_{sorted}\leftarrow L_{sorted}+a$triangleright Add to sorted list
@@
14:      $L_{circular}\leftarrow DFS(L_{i},dep)$triangleright Circular dependencies
15:      $L_{sorted}\leftarrow L_{sorted}+L_{circular}$triangleright Add a group to sorted list
@@ C.1 Symbolic Execution-Based Equivalence
Symbolic execution explores all potential execution paths of a program by using symbolic inputs to generate constraints [King, 1976, Baldoni et al., 2018, Coward, 1988]. While theoretically powerful, this method is impractical for verifying C-to-Rust equivalence due to differences in language features. For instance, Rust’s RAII (Resource Acquisition Is Initialization) pattern automatically inserts destructors for memory management, while C relies on explicit malloc and free calls. These differences cause mismatches in compiled code, making it difficult for symbolic execution engines to prove equivalence. Additionally, Rust’s compiler adds safety checks (e.g., array boundary checks), which further complicate equivalence verification.
@@ C.2 Fuzz Testing-Based Equivalence
Fuzz testing generates random or mutated inputs to test whether program outputs match expected results [Zhu et al., 2022, Miller et al., 1990, Liang et al., 2018]. While more practical than symbolic execution, fuzz testing faces challenges in constructing meaningful inputs for real-world programs. For example, testing a URL parsing function requires generating valid URLs with specific formats, which is non-trivial. For large C programs, this difficulty scales, making it infeasible to produce high-quality test cases for every translated Rust function.
Appendix D Dataset Details
D.1 TransCoder-IR Dataset [Szafraniec et al., 2023]
@@ Appendix D Dataset Details
D.2 Project CodeNet [Puri et al., 2021]
@@
D.3 Real-World Projects
For § 6.4, we use two real-world projects for evaluation: avl-tree [^4] https://github.com/xieqing/avl-tree and urlparser [^5] https://github.com/jwerle/url.h. Both projects are written in C and have some non-trivial C code features. While earlier works may have used different versions, we selected these projects because they include end-to-end tests, enabling us to evaluate the correctness of the translated code. To make the projects fit the input requirements, we use the cpp preprocessor to expand macros and combine the entire project into a single file. The avl-tree project contains 12 end-to-end tests to test the different functionalities of the AVL tree implementation. The urlparser project contains 3 end-to-end tests to test the different functionalities of the URL, we manually create 7 additional end-to-end tests to test the different functionalities of the URL parser.
Appendix E LLM Configurations
Table 7 shows our configurations for different LLMs in evaluation. All other hyper-parameters, like Top-P or Top-K, are set as the model’s default values.

### v3

@@ A.2 Tabular Summary
Table 3: Key Differences Between C and Rust
Appendix B Preprocessing and Task Division
B.1 Preprocessing of C Files
To support real-world C projects, SACTOR parses the compile commands generated by the make tool, extracting relevant flags for preprocessing, parsing, compilation, linking, and third-party tools’ use.
C source files usually contain preprocessing directives, such as #include, #define, #ifdef, #endif, etc., which we need to resolve before parsing C files. For #include, we copy and expand non-system headers recursively while keeping #include of system headers intact, because included non-system headers contain project-specific definitions such as structs and enums that the LLM has not known while system headers’ contents are known to the LLM and expanding them would unnecessarily introduce too much noise. For other directives, we pass relevant C project compile flags to the C preprocessor from GCC to resolve them.
B.2 Algorithm for Task Division
@@
[…]
@@
9:   $L_{sorted}\leftarrow L_{sorted}+a$ $\triangleright$ Add to sorted list
@@
14:  $L_{circular}\leftarrow DFS(L_{i},dep)$ $\triangleright$ Circular dependencies
15:  $L_{sorted}\leftarrow L_{sorted}+L_{circular}$ $\triangleright$ Add a group to sorted list
@@ C.1 Symbolic Execution-Based Equivalence
Symbolic execution explores all potential execution paths of a program by using symbolic inputs to generate constraints [king1976symbolic, baldoni2018survey, coward1988symbolic]. While theoretically powerful, this method is impractical for verifying C-to-Rust equivalence due to differences in language features. For instance, Rust’s RAII (Resource Acquisition Is Initialization) pattern automatically inserts destructors for memory management, while C relies on explicit malloc and free calls. These differences cause mismatches in compiled code, making it difficult for symbolic execution engines to prove equivalence. Additionally, Rust’s compiler adds safety checks (e.g., array boundary checks), which further complicate equivalence verification.
@@ C.2 Fuzz Testing-Based Equivalence
Fuzz testing generates random or mutated inputs to test whether program outputs match expected results [zhu2022fuzzing, miller1990empirical, liang2018fuzzing]. While more practical than symbolic execution, fuzz testing faces challenges in constructing meaningful inputs for real-world programs. For example, testing a URL parsing function requires generating valid URLs with specific formats, which is non-trivial. For large C programs, this difficulty scales, making it infeasible to produce high-quality test cases for every translated Rust function.
Appendix D An Example of the Test Harness
Here, we provide an example of the test harness used to verify the correctness of the translated code in Figure 4, which is used to verify the idiomatic Rust code. In this example, the concat_str_idiomatic function is the idiomatic translation we are testing, while the concat_str_c function is the test harness function that can be linked back to the original C code. where a string and an integer are passed as input, and an owned string is returned. Input strings are converted from C’s char* to Rust’s &str, and output strings are converted from Rust’s String back to C’s char*.
Figure 4: Test harness used for verifying concat_str translation
Appendix E An Example of SACTOR Translation Process
To demonstrate the translation process of SACTOR, we present a straightforward example of translating a C function to Rust. The C program includes an atoi function that converts a string to an integer, and a main function that parses command-line arguments and calls the atoi function. The C code is shown in Figure 5(a).
(a) C implementation of atoi
(b) Unidiomatic Rust translation from C
(c) Idiomatic Rust translation from unidiomatic Rust
Figure 5: SACTOR translation process for atoi program
We assume that there are numerous end-to-end tests for the C code, allowing SACTOR to use them for verifying the correctness of the translated Rust code.
First, the divider will divide the C code into two parts: the atoi function and the main function, and determine the translation order is first atoi and then main, as atoi is the dependency of main and the atoi function is a pure function.
Next, SACTOR proceeds with the unidiomatic translation, converting both functions into unidiomatic Rust code. This generated code will keep the semantics of the original C code while using Rust syntax. Once the translation is complete, the unidiomatic verifier executes the end-to-end tests to ensure the correctness of the translated function. If the verifier passes all tests, SACTOR considers the unidiomatic translation accurate and progresses to the next function. If any test fails, SACTOR will retry the translation process using the feedback information collected from the verifier, as described in § 4.3. After translating all sections of the C code, SACTOR will combine the unidiomatic Rust code segments to form the final unidiomatic Rust code. The unidiomatic Rust code is shown in Figure 5(b).
Then, the SACTOR will start the idiomatic translation process and translate the unidiomatic Rust code into idiomatic Rust code. The idiomatic translator requests the LLM to adapt the C semantics into idiomatic Rust, eliminating any unsafe and non-idiomatic constructs, as detailed in § 4.2. Based on the same order, the SACTOR will translate two functions accordingly, and using the idiomatic verifier to verify and provide the feedback to the LLM if the verification fails. After all parts of the Rust code are translated into idiomatic Rust, verified, and combined, the SACTOR will produces the final idiomatic Rust code. The idiomatic Rust code is shown in Figure 5(c), representing the final output of SACTOR.
Appendix F Dataset Details
Dataset Size Preprocessing E2E Tests Coverage (Line/Func)
TransCoder-IR [transcoderir] 100 Removed buggy programs (compilation/memory errors) and entries with existing Rust Present 97.97% / 99.5%
Project CodeNet [codenet] 100 Filtered for external-input programs (argc/argv); auto-generated tests Generated 94.37% / 100%
CRust-Bench [khatry2025crust] 50 Excluded unsupported patterns; combine code of each sample to a single lib.c Present 76.18% / 80.98%
libogg [libogg] 1 None. Each component of the library is contained within a single C file. Present 83.3% / 75.3%
Table 4: Summary of datasets and real-world code-bases used for evaluation; coverage audited with gcov on the tests exercised in our pipeline.
F.1 TransCoder-IR Dataset [transcoderir]
@@
F.2 Project CodeNet [codenet]
@@
F.3 CRust-Bench [khatry2025crust]
CRust-Bench is a repository-level benchmark for C-to-safe-Rust transpilation. It collects 100 real-world C repositories (the CBench suite) and pairs each with a manually written, safe Rust interface and a set of tests that assert functional correctness. By evaluating full repositories rather than isolated functions, CRust-Bench surfaces challenges common in practice, such as complex, pointer-rich APIs. In our evaluation, we use a 50-sample subset in CRust-Bench, which exclude entries that are out of scope for our pipeline (e.g., circular type or function dependencies and compiler-specific intrinsics that do not map cleanly). For each selected sample, we reuse the upstream end-to-end tests and relink them so that calls exercise our translated code; build environments and link flags follow the sample’s configuration.
F.4 libogg [libogg]
libogg is the reference implementation of the Ogg multimedia container. Ogg is a stream-oriented format that frames, timestamps, and multiplexes compressed media bitstreams (e.g., audio/video) into a robust, seekable stream. The libogg distribution contains only the Ogg container library (codecs such as Vorbis or Theora are hosted separately). In our case study, the codebase comprises roughly 2,041 lines of code (excluding tests), six struct definitions, three global variables, and 77 exported functions. We use the project’s upstream tests and build scripts. This single-project evaluation complements the CRust-Bench subset by focusing on non-trivial structs, buffers, and pointer manipulation in a real-world C library.
Appendix G LLM Configurations
Table 5 shows our configurations for different LLMs in evaluation. All other hyperparameters (e.g., Top-P, Top-K) use provider defaults. As GPT-5 does not support temperature setting, we use its default temperature.
Model Version Temperature
GPT-4o gpt-4o-2024-08-06 0
Claude 3.5 Sonnet claude-3-5-sonnet-20241022 0
Gemini 2.0 Flash gemini-2.0-flash-exp 0
Llama 3.3 Instruct 70B Llama 3.3 Instruct 70B1 0
DeepSeek-R1 DeepSeek-R1 671B2 0
GPT-5 gpt-5-2025-08-07 default
  • 1
    https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
  • 2
    https://huggingface.co/deepseek-ai/DeepSeek-R1